Suppose that 10% of all steel shafts produced by a certain process are nonconforming but can be re- worked (rather than having to be scrapped). Consider a random sample of 200 shafts, and let X denote the number among these that are nonconforming and can be reworked. What is the (approximate) probability that X is a) At most 30? [2] b) Between 15 and 25 (both inclusive)? [2] c) Assume that the probability of at most x shafts being nonconforming is 0.9906. Find x. [3]2. Suppose that 10% of all steel shafts produced by a certain process are nonconforming but can be re- worked (rather than haI really want to know the answer for Part c
Answer #1

As we are talking about approximate probabilities here, we would be using the normal approximation to binomial distribution here to get the required probabilities here.

The distribution here is given as:

X \sim Bin(n = 200, p = 0.1)

X \sim N(\mu = np, \sigma^2 = np(1-p) )

X \sim N(\mu = 200*0.1 , \sigma^2 = 200*0.1*0.9)

X \sim N(\mu = 20 , \sigma^2 = 18)

a) The probability here is computed as:
P(X <= 30)

Applying the continuity correction, we have here:
P(X < 30.5)

Converting it to a standard normal variable, we have here:

30.5 - 20 P(Z < /18

P(Z < 2.48)

Getting it from the standard normal tables, we have here:

P(Z < 2.48) = 0.9933

Therefore 0.9933 is the required probability here.

b) The probability here is computed as:
P(15 <= X <= 25)

Applying the continuity correction, we have here:

P( 14.5 < X < 25.5)

Converting it to a standard normal variable, we have here:

P( \frac{14.5 - 20}{\sqrt{18}} <Z < \frac{25.5 - 20}{\sqrt{18}})

P( - 1.2964 <Z < 1.2964)

= P( Z < 1.2964) - P(Z < - 1.2964 )

Getting it from the standard normal tables, we have here:

= 0.9026 - (1 - 0.9026) = 0.8052

Therefore 0.8052 is the required probability here.

c) From standard normal tables, we have here:
P(Z < 2.349) = 0.9906

Therefore, the value of x here is computed as:

\frac{x - 20}{\sqrt{18}} = 2.349

x = 20 + 2.349\sqrt{18} = 29.97

Therefore 30 is the required value of x here.